# One Communication Round is All It Needs for Federated Fine-Tuning Foundation Models

### v1

@@ Abstract
The recent advancement of foundation models (FMs) has increased the demand for fine-tuning these models on large-scale cross-domain datasets. To address this, federated fine-tuning has emerged as a solution, allowing FMs to be fine-tuned on distributed datasets across multiple devices while ensuring data privacy. However, the substantial parameter size of FMs and the multi-round communication required by federated fine-tuning algorithms result in prohibitively high communication costs, challenging the practicality of federated fine-tuning. In this paper, we are the first to reveal, both theoretically and empirically, that the traditional multi-round aggregation algorithms may not be necessary for federated fine-tuning large FMs. Our experiments reveal that a single round of aggregation (i.e., one-shot federated fine-tuning) yields a global model performance comparable to that achieved through multiple rounds of aggregation. Through rigorous mathematical and empirical analyses, we demonstrate that large FMs, due to their extensive parameter sizes and pre-training on general tasks, achieve significantly lower training loss in one-shot federated fine-tuning compared to smaller models. Our extensive experiments show that one-shot federated fine-tuning not only reduces communication costs but also enables asynchronous aggregation, enhances privacy, and maintains performance consistency with multi-round federated fine-tuning on both text generation and text-to-image generation tasks. Our findings have the potential to revolutionize federated fine-tuning in practice, enhancing efficiency, reducing costs, and expanding and accessibility for FMs.
One Communication Round is All It Needs for Federated Fine-Tuning Foundation Models
Ziyao Wang1, Bowei Tian1, Yexiao He1, Zheyu Shen1, Luyang Liu2, Ang Li1 1. University of Maryland, College Park 2. Google DeepMind {ziyaow,boweitian,yexiaohe,zyshen,angliece}@umd.edu, luyangliu@google.com
1 Introduction
@@
Cutting-edge foundation models (FMs) demonstrate remarkable versatility across various domains. Notably, large language models (LLMs) like GPT-4 Achiam et al. (2023), Gemma Team et al. (2024), and Llama Touvron et al. (2023b) excel in tasks such as translation, question answering (QA), chat assistant, and math. Similarly, stable diffusion models can generate diverse images based on textual descriptions. Achieving such versatility requires fine-tuning these FMs on cross-domain datasets. However, this process faces significant challenges in real-world scenarios due to the valuable datasets residing on devices owned by organizations or individuals, raising privacy concerns. To address these privacy issues, researchers have proposed using federated learning (FL) Zhang et al. (2021) for distributed fine-tuning of FMs, a process known as federated fine-tuning. Federated fine-tuning allows distributed clients to collaboratively fine-tune a global FM on specific tasks without disclosing their private data.
@@
Traditional FL requires multiple communication rounds between clients and the server to ensure the global model convergence McMahan et al. (2017). However, the substantial parameter size of FMs (typically in billions) results in significant communication overhead. Many devices lack the capability to repeatedly communicate model parameters of this scale. While previous works adopt parameter-efficient fine-tuning (PEFT) methods such as low-rank adaptation (LoRA) Hu et al. (2021) to reduce the number of trainable and communicated parameters, the high communication requirements of federated fine-tuning remain a practical limitation.
@@
Figure 1: The distinct performances of one-shot federated learning between small models and large FMs. The horizontal axis represents multi-round FL accuracy, while the vertical axis represents one-shot FL accuracy. The ResNet-18 and LSTM are trained and tested on CIFAR-10 and Shakespeare respectively. Other models are fine-tuned on Wizard dataset and tested on ARC Easy. The closer points are to the dashed line means the closer accuracy between one-shot and multi-round FL.
@@
Unexpectedly, our recent experiments have discovered an emergent capability of FMs that could fundamentally shift the approach to federated fine-tuning. We find that with sufficient local fine-tuning epochs, a single communication round is all it needs to effectively fine-tune FMs, which is called one-shot federated fine-tuning. Figure 1 highlights the performance comparisons between one-shot FL and traditional multi-round FL, maintaining the same total number of local epochs. While one-shot FL underperforms multi-round FL for smaller models (e.g., ResNet-18 and LSTM), it achieves comparable performance for larger FMs (e.g., GPT-2, Llama, etc). This unique discovery challenges the conventional belief that multiple communication rounds are essential for the federated fine-tuning of FMs. Instead, we demonstrate that FMs can achieve convergence with just a single aggregation of well-fine-tuned local models. This paper explores this innovative finding, providing rigorous theoretical analysis and compelling empirical evidence to validate the effectiveness of one-shot FL for federated fine-tuning FMs.
@@
The introduction of one-shot FL brings transformative benefits. First, it dramatically reduces communication costs. One-Shot FL slashes communication overhead by a factor of $\frac{1}{T}$, where $T$ represents the number of communication rounds in traditional federated fine-tuning. This reduction is a game-changer for devices with limited bandwidth. Second, one-shot FL enables seamless asynchronous training. This flexibility removes the bottleneck of server waiting times, ensuring uninterrupted training regardless of client connectivity or resource limitations. The process becomes far more robust and efficient. Third, one-shot FL offers enhanced security against prevalent client-side federated learning attacks. Attacks like client-side model inversion and gradient inversion, which depend on multiple global model updates, are rendered ineffective. This significantly bolsters the integrity of the training process.
@@
  • Novel Discovery: To the best of our knowledge, we are the first to discover that one-round aggregation is sufficient for federated fine-tuning large FMs.
  • Theoretical Analysis: We theoretically demonstrate the relationship between the error of one-shot federated fine-tuning and model smoothness, fine-tuning model update, number of fine-tuning rounds, and the second norm of original model parameters. Our analysis, supported by experiments, reveals that large FMs are smoother, exhibit smaller model updates, and require fewer fine-tuning epochs than smaller models, resulting in significantly lower one-shot federated fine-tuning error.
  • Experimental Validation: We conduct extensive experiments on six FMs and three tasks, demonstrating that one-shot federated fine-tuning achieves performance comparable to multi-round federated fine-tuning, particularly for models with over 1 billion parameters. Experimental results also surprisingly show that LoRA outperforms full fine-tuning in the context of one-shot federated fine-tuning.
2 Preliminary
Federated Learning Paradigm of Small Models. In FL, the primary objective is to optimize a global objective function $F(\bm{w})$, which is weighted average of the local objective functions from $m$ clients Wang et al. (2020b):
$F(\bm{w})=\sum_{i=1}^{m}p_{i}F_{i}(\bm{w})$ (1)
where $\bm{w}$ represents the model parameters and $p_{i}$ is the scaling factor for averaging. To protect the data privacy of each client, the server cannot access the local dataset. Thus, the local objective function $F_{i}(\bm{w})$ remains unknown to the server. FedAvg McMahan et al. (2017) algorithm provides a distributed training algorithm to facilitate privacy-conscious training. It allows multiple clients to train the model on their local datasets and aggregates locally trained models on the server at the end of each communication round. In $t$-th communication round, the global model update rule of FedAvg is:
$\bm{w}^{(t+1,0)}-\bm{w}^{(t,0)}=\alpha^{(t)}\sum_{i=1}^{m}p_{i}\Delta_{i}^{(t)},t\in[0,T-1]$ (2)
where $\bm{w}^{(t,0)}$ is the model weights in $t$-th communication round and 0-th local epoch, which represents the global model in $t$-th round. $T$ is the total number of communication rounds, $\alpha^{(t)}$ is the global learning rate, and $\Delta_{i}^{(t)}$ is the local model update in $t$-th round. $\Delta_{i}^{(t)}$ is the accumulative model update of $k$ local stochastic gradient descent (SGD) steps:
$\Delta_{i}^{(t)}=\sum_{j=1}^{k}\beta_{i}^{(t,j)}g_{i}(\bm{w}_{i}^{(t,j)})$ (3)
@@ 2 Preliminary
Local datasets in FL are typically heterogeneous, leading to differences in local objectives. Therefore, FL usually converges more slowly than centralized machine learning. This slow convergence necessitates a large number of global communication rounds and local epochs to achieve satisfactory performance. For example, experiment results in Reddi et al. (2020) show that the ResNet-18 model requires more than 2000 and 4000 communication rounds to converge on CIFAR-10 Krizhevsky et al. (2009) and CIFAR-100 respectively. Even for simple natural language processing tasks such as Shakespeare, an RNN model needs more than 50 rounds to converge. The requirement for multi-round communication rounds introduces several significant drawbacks. First, clients must frequently exchange model parameters with the server, which can be prohibitively expensive in certain constrained scenarios or on devices with limited resources. Second, repeated invocation of computational resources for training increases the overall computational overhead. Additionally, the multi-round communication approach leads to excessive energy consumption, synchronization difficulties, and challenges in maintaining privacy protection. Thus, optimizing FL algorithms to minimize the number of communication rounds is an essential research direction in FL.
@@
Federated Fine-Tuning Foundation Models. Foundation models (FMs) Zhou et al. (2023) refer to pre-trained deep learning models with a vast number of parameters, typically in the order of billions. These FMs are trained on broad data at scale and are adaptable to a wide range of downstream tasks when fine-tuned on domain-specific datasets Bommasani et al. (2021). Since domain-specific datasets are often distributed across multiple devices, FL offers an important paradigm for fine-tuning FMs while preserving data privacy.
@@
Federated fine-tuning adopts the same FedAvg algorithm in Eq. 1 and Eq. 2 to aggregate the local model updates. The key difference lies in the model parameter size. The parameter size of large FMs is usually hundreds of times greater than that of small models, resulting in a significant increase in the computation resources and communication overhead required for federated fine-tuning. Given the network communication capabilities of commonly used devices, performing multi-round synchronized communication of large model parameters between servers and clients is virtually impossible. Although parameter-efficient fine-tuning algorithms like LoRA Hu et al. (2021) have been adopted, the communication overhead remains excessively high, hindering practical application.
@@
One-Shot Federated Learning. To reduce communication overhead in FL, recent works have focused on one-shot FL Jhunjhunwala et al. (2024); Guha et al. (2019); Gong et al. (2021); Li et al. (2020); Zhou et al. (2020); Yang et al. (2024), which uses a single communication round to obtain the global model. These algorithms often employ knowledge distillation or neuron-matching methods to optimize the global model. However, these approaches require additional data or computation. Knowledge distillation often necessitates auxiliary public datasets or external generative models, and neuron matching requires additional computation on both clients and the server. Despite these additional resource requirements, the performance of one-shot FL has historically been inferior to standard multi-round FL. For instance, experiments in Jhunjhunwala et al. (2024) show that one-shot FL achieves only 50% accuracy on the CIFAR-10 dataset, which is 20% lower than the accuracy achieved with 5-round FL.
@@
However, our recent experiments have uncovered greater potential for one-shot federated fine-tuning large FMs. As shown in Figure 1, one-shot FL for large models does not show a significant performance gap compared to multi-round FL, which is commonly observed with smaller models. In fact, when the total number of local epochs is the same, the performance of large models fine-tuned by one-shot FL is comparable to that of multi-round FL. Additionally, in fine-tuning larger models such as Llama-13b, one-shot FL even performed slightly better than multi-round FL. These results, along with the experiment results in Section 4, suggest that traditional multi-round FL algorithms may no longer be necessary for federated fine-tuning large FMs. Large FMs can effectively learn downstream tasks from distributed clients with just a single communication round, opening up new possibilities for federated fine-tuning applications.
@@
Although we have observed consistently good performance with one-shot federated fine-tuning, the reasons for the divergent results between one-shot FL in fine-tuning large models and training small models remain unexplored. In the next section, we will delve into this phenomenon through theoretical analysis.
3 Theoretical Analysis of One-Shot Federated Fine-Tuning
@@
For a multi-round FL algorithm, if the total number of communication rounds is $T$ and the number of local epochs for each round is $k$, according to Eq. 2 the global model parameters after FL satisfy:
$\bm{w}^{(T,0)}-\bm{w}^{(0,0)}=\sum_{t=1}^{T}\alpha^{(t)}\sum_{i=1}^{m}p_{i}\Delta_{i}^{(t)},$ (4)
@@
$\Delta_{i}=\sum_{t=1}^{T}\Delta_{i}^{(t)}=\sum_{t=1}^{T}\sum_{j=1}^{k}\beta_{i}^{(t,j)}g_{i}(\bm{w}_{i}^{(t,j)}),$ (5)
@@
$\Delta_{i}=\sum_{j=1}^{Tk}\beta_{i}^{(0,j)}g_{i}(\bm{w}_{i}^{(0,j)}),$ (6)
Here we set the number of epochs per client to $Tk$ since we are trying to match the total number of epochs with the multi-round FL. The reason why the one-shot FL performs worse than the multi-round FL in small models lies in the difference between the local model updates in Eq. 5 and Eq. 6. In Eq. 5, after the $t$-th communication round, the local training starts from the updated global model parameter $\bm{w}^{(t,0)}$ sent by the server, which is aggregated from all the local model updates in $t$-th round and contains richer local knowledge. Therefore, the client can compute a more accurate gradient $g_{i}(\bm{w}_{i}^{(t,j)})$ based on the updated model. On the contrary, in one-shot FL (Eq. 6), even with the same total local epochs, clients can only continuously train the local models without global information. The poor performance of one-shot FL is due to the gradients calculated on the local models being less accurate than those calculated on the aggregated global model. This local error can be expressed in mathematical form:
$\varepsilon_{i}=\sum_{j=k+1}^{Tk}\beta_{i}^{(0,j)}[(g_{i}(\bm{w}_{i}^{(0,j)})-g_{i}(\bm{w}_{i}^{(t,j-kt)})],$ (7)
[…]
$\varepsilon\leq\sum_{i=1}^{m}\varepsilon_{i},$ (8)
The global error can be further simplified by the following assumptions.
@@
[…]
@@
Theorem 1 (The one-shot FL error is related to $L$, $\tau$, epoch numbers $Tk$, and number of clients $m$). Under Assumptions 1 and 2, ignoring the difference of learning rates in one-shot and multi-round FL and the difference of client numbers (i.e., set client number $m$ to 1), the error of one-shot FL $\varepsilon$ can be bounded as follows:
$\|\varepsilon\|\leq\Gamma\|\bm{w}^{(0,0)}\|,\ \text{where}\ \Gamma=L\tau Tk$ (9)
@@
This equation indicates that with lower values of $L$, $\tau$, $T$, $k$, and $m$, the model update of one-shot FL will be closer to that of multi-round FL. Conversely, if a neural network has a highly complex loss landscape, large training dynamics, or requires a large number of rounds to converge, the error $\varepsilon$ will be large, leading to poor performance of one-shot FL. Since our experiments have shown that LLMs exhibit significant advantages over small models in one-shot learning, we conduct experiments on the factors in Equation 9 to provide a detailed explanation of this phenomenon.

### v2

# Revisiting Federated Fine-Tuning: A Single Communication Round is Enough for Foundation Models

@@ Abstract
The recent advancement of foundation models (FMs) has increased the demand for fine-tuning these models on large-scale cross-domain datasets. To address this, federated fine-tuning has emerged, allowing FMs to be fine-tuned on distributed datasets across multiple devices while ensuring data privacy. However, the substantial parameter size and the multi-round communication in federated learning algorithms result in prohibitively high communication costs, challenging the practicality of federated fine-tuning. In this paper, we identify and analyze, both theoretically and empirically, that the traditional multi-round aggregation algorithms may not be necessary for federated fine-tuning large FMs. Our experiments reveal that a single round of aggregation (i.e., one-shot federated fine-tuning) yields a global model performance comparable to that achieved through multiple rounds of aggregation. Through rigorous mathematical and empirical analyses, we demonstrate that large FMs, due to their extensive parameter sizes and pre-training on general tasks, achieve significantly lower training loss in one-shot federated fine-tuning compared to smaller models. Our extensive experiments show that one-shot federated fine-tuning significantly reduces communication costs. It also has the potential to enable asynchronous aggregation, enhances privacy, and maintains performance consistency with multi-round federated fine-tuning on both text generation and text-to-image generation tasks. Our findings provide insights to revolutionize federated fine-tuning in practice, enhancing efficiency, reducing costs, and expanding accessibility for FMs.
I Introduction
@@
Cutting-edge foundation models (FMs) demonstrate remarkable versatility across various domains. Notably, large language models (LLMs) like GPT-4 [achiam2023gpt], Gemma [team2024gemma], and Llama [touvron2023llama2] excel in tasks such as translation, question answering (QA), chat assistant, and math. Similarly, diffusion models [song2020score, yang2023diffusion] can generate diverse images based on textual descriptions. Achieving such versatility requires fine-tuning these FMs on cross-domain datasets. However, this process faces significant challenges in real-world scenarios due to the valuable datasets residing on devices owned by organizations or individuals, raising privacy concerns. To address these privacy issues, researchers have proposed using federated learning (FL) [mcmahan2017communication, li2020lotteryfl, li2021hermes] for distributed fine-tuning of FMs, a process known as federated fine-tuning [zhang2024towards, ye2024openfedllm, yao2409federated]. Federated fine-tuning allows distributed clients to collaboratively fine-tune a global FM on specific tasks without disclosing their private data.
@@
Traditional FL requires multiple communication rounds between clients and the server to ensure the global model convergence [mcmahan2017communication]. However, the substantial parameter size of FMs (typically in billions) results in significant communication overhead [ghiasvand2024communication]. Many devices lack the capability to repeatedly communicate model parameters of this scale [liu2022chronos]. While previous works adopt parameter-efficient fine-tuning (PEFT) methods such as low-rank adaptation (LoRA) [hu2021lora, yang2024low, shen2025edgelora, wang2025prada] to reduce the number of trainable and communicated parameters, the high communication requirements of federated fine-tuning remain a practical limitation [sun2024improving, wang2024flora, yi2023pfedlora, wang2025fedmabench].
@@
Figure 1: The distinct performances of one-shot federated learning between small models and large FMs. The horizontal axis represents multi-round FL accuracy, while the vertical axis represents one-shot FL accuracy. The ResNet-18 and LSTM are trained and tested on CIFAR-10 and Shakespeare respectively. Other models are fine-tuned on Wizard dataset and tested on ARC Easy. The closer the points are to the dashed line means the accuracy of one-shot and multi-round FL are closer in the corresponding model.
@@
Unexpectedly, our recent experiments have discovered an emergent capability of FMs that could fundamentally shift the approach to federated fine-tuning. We find that with sufficient local fine-tuning epochs, a single communication round is enough to effectively fine-tune FMs, which is called one-shot federated fine-tuning [guha2019one]. Fig. 1 highlights the performance comparisons between one-shot FL and traditional multi-round FL, maintaining the same total number of local epochs. While one-shot FL underperforms multi-round FL for smaller models (e.g., ResNet-18 and LSTM), it achieves comparable performance for larger FMs (e.g., GPT-2, Llama, etc). This unique discovery challenges the conventional belief that multiple communication rounds are essential for the federated fine-tuning of FMs. Instead, we demonstrate that FMs can achieve convergence with just a single aggregation of well-fine-tuned local models. This paper explores this innovative finding, providing rigorous theoretical analysis and compelling empirical evidence to validate the effectiveness of one-shot federated fine-tuning FMs. Our theory for why one-shot FL works well in larger models also sheds light on the success of task vectors [hendel2023context, wang2024moderator], task arithmetic [ilharco2022editing], and model merging [zhao2024texttt] in FMs.
@@
Beyond the theoretical contributions, an effective one-shot FL also offers advantages for the practical deployment of FL systems. First, it dramatically reduces communication costs. One-Shot FL slashes communication overhead by a factor of $\frac{1}{T}$, where $T$ represents the number of communication rounds in traditional federated fine-tuning. This reduction is a game-changer for devices with limited bandwidth. Second, one-shot FL enables seamless asynchronous training. This flexibility removes the bottleneck of server waiting times, ensuring uninterrupted training regardless of client connectivity or resource limitations. The process becomes far more robust and efficient. Third, one-shot FL offers enhanced security against prevalent client-side federated learning attacks. Attacks like client-side model inversion and gradient inversion, which depend on multiple global model updates, are rendered ineffective. This significantly bolsters the integrity of the training process.
@@
  • Empirical Insight: We identify and systematically validate the surprising effectiveness of one-shot federated fine-tuning of large foundation models.
  • Theoretical Analysis: We theoretically demonstrate the relationship between the error of one-shot federated fine-tuning and model smoothness, fine-tuning model update, and number of fine-tuning rounds. Our analysis, supported by experiments, reveals that large FMs are smoother, exhibit smaller model updates, and require fewer fine-tuning epochs than smaller models, resulting in significantly lower one-shot federated fine-tuning errors.
  • Experimental Validation: We conduct extensive experiments on six FMs and three tasks, demonstrating that one-shot federated fine-tuning achieves performance comparable to multi-round federated fine-tuning, particularly for models with over 1 billion parameters. Experimental results also surprisingly show that LoRA outperforms full fine-tuning in some cases of one-shot federated fine-tuning.
II Preliminary
Federated Learning Paradigm of Small Models. In FL, the primary objective is to optimize a global objective function $F(\bm{w})$, which is weighted average of the local objective functions from $m$ clients [wang2020tackling]:
$F(\bm{w})=\sum_{i=1}^{m}p_{i}F_{i}(\bm{w}),$ (1)
where $\bm{w}$ represents the model parameters and $p_{i}$ is the scaling factor. To protect the data privacy of each client, the server cannot access the local dataset. Thus, the local objective function $F_{i}(\bm{w})$ remains unknown to the server. FedAvg [mcmahan2017communication] algorithm provides a distributed training algorithm to facilitate privacy-conscious training. It allows multiple clients to train the model on their local datasets and aggregates locally trained models on the server at the end of each communication round. In the $t$-th communication round, the global model update rule of FedAvg is:
$\bm{w}^{(t+1,0)}-\bm{w}^{(t,0)}=\alpha^{(t)}\sum_{i=1}^{m}p_{i}\Delta_{i}^{(t)},\quad t\in[0,T-1],$ (2)
where $\bm{w}^{(t,0)}$ is the model weights in $t$-th round and 0-th local epoch, which represents the global model in $t$-th round. $T$ is the total number of communication rounds, $\alpha^{(t)}$ is the global learning rate, and $\Delta_{i}^{(t)}$ is the local model update in $t$-th round. $\Delta_{i}^{(t)}$ is the accumulative model update of $k$ local stochastic gradient descent (SGD) steps:
$\Delta_{i}^{(t)}=\sum_{j=1}^{k}\beta_{i}^{(t,j)}g_{i}(\bm{w}_{i}^{(t,j)}),$ (3)
@@ II Preliminary
Local datasets in FL are typically heterogeneous, leading to differences in local objectives. Therefore, FL usually converges more slowly than centralized machine learning [mcmahan2017communication, wang2023fedhyper]. This slow convergence necessitates a large number of global communication rounds and local epochs to achieve satisfactory performance. For example, experimental results in [reddi2020adaptive] show that the ResNet-18 model requires more than 2000 and 4000 communication rounds to converge on CIFAR-10 [krizhevsky2009learning] and CIFAR-100 respectively. Even for simple natural language processing tasks such as Shakespeare, an RNN model needs more than 50 rounds to converge. The requirement for multi-round communication rounds introduces several significant drawbacks. First, clients must frequently exchange model parameters with the server, which can be prohibitively expensive in certain constrained scenarios or on devices with limited resources. Second, repeated invocation of computational resources for training increases the overall computational overhead. Additionally, the multi-round communication approach leads to excessive energy consumption, synchronization difficulties, and challenges in maintaining privacy protection. Thus, optimizing FL algorithms to minimize the number of communication rounds is an essential research direction in FL.
@@
Federated Fine-Tuning Foundation Models. Foundation models [zhou2023comprehensive] refer to pre-trained deep learning models with a vast number of parameters, typically in the order of billions. These FMs are trained on broad data at scale and are adaptable to a wide range of downstream tasks when fine-tuned on domain-specific datasets [bommasani2021opportunities]. Since domain-specific datasets are often distributed across multiple devices, FL offers an important paradigm for fine-tuning FMs while preserving data privacy.
@@
Federated fine-tuning [sun2024improving, wang2024flora, cheng2021fine, woisetschlager2024federated, hilmkil2021scaling] adopts the same FedAvg algorithm in Eq. 1 and Eq. 2 to aggregate the local model updates. The key difference lies in the model parameter size. The parameter size of FMs is usually hundreds of times greater than that of small models, resulting in a significant increase in the computation resources and communication overhead required for federated fine-tuning. Given the network communication capabilities of commonly used devices, performing multi-round synchronized communication of large model parameters between servers and clients is virtually impossible. Although parameter-efficient fine-tuning algorithms like LoRA [hu2021lora] have been adopted in federated fine-tuning, the communication overhead remains excessively high, hindering practical application.
@@
One-Shot Federated Learning. To reduce communication overhead in FL, recent works have focused on one-shot FL [jhunjhunwala2024fedfisher, guha2019one, gong2021ensemble, li2020practical, zhou2020distilled, yang2024exploring], which uses a single communication round to obtain the global model. These algorithms often employ knowledge distillation or neuron-matching methods to optimize the global model. However, these approaches require additional data or computation. Knowledge distillation often necessitates auxiliary public datasets or external generative models, and neuron matching requires additional computation on both clients and the server. Despite these additional resource requirements, the performance of one-shot FL has historically been inferior to standard multi-round FL. For instance, experiments in [jhunjhunwala2024fedfisher] show that one-shot FL achieves only 50% accuracy on the CIFAR-10 dataset, which is 20% lower than the accuracy achieved with 5-round FL.
@@
However, our recent experiments have uncovered greater potential for one-shot federated fine-tuning large FMs. As shown in Fig. 1, one-shot FL for large models does not show a significant performance gap compared to multi-round FL, which is commonly observed with smaller models. In fact, when the total number of local epochs is the same, the performance of large models fine-tuned by one-shot FL is comparable to that of multi-round FL. Additionally, in fine-tuning larger models such as Llama-13b, one-shot FL even performed slightly better than multi-round FL. These results, along with the experiment results in Section IV, suggest that traditional multi-round FL algorithms may no longer be necessary for federated fine-tuning large FMs. Large FMs can effectively learn downstream tasks from distributed clients with just a single communication round, opening up new possibilities for federated fine-tuning applications.
@@
Although we have observed consistently good performance with one-shot federated fine-tuning, the reasons behind this phenomenon remain unexplored. In the next section, we will delve into this phenomenon through theoretical analysis.
III Theoretical Analysis of One-Shot Federated Fine-Tuning
@@
For a multi-round FL algorithm, if the total number of communication rounds is $T$ and the number of local SGD steps for each round is $k$, according to Eq. 2 the global model parameters after FL satisfy:
$\bm{w}^{(T,0)}-\bm{w}^{(0,0)}=\sum_{t=0}^{T-1}\alpha^{(t)}\sum_{i=1}^{m}p_{i}\Delta_{i}^{(t)},$ (4)
@@
$\Delta_{i}=\sum_{t=0}^{T-1}\Delta_{i}^{(t)}=\sum_{t=0}^{T-1}\sum_{j=1}^{k}\beta_{i}^{(t,j)}g_{i}(\bm{w}_{i}^{(t,j)}).$ (5)
@@
$\Delta_{i}=\sum_{j=1}^{Tk}\beta_{i}^{(0,j)}g_{i}(\bm{w}_{i}^{(0,j)}).$ (6)
Here we set the number of steps per client to $Tk$ since we are trying to match the total number of steps with the multi-round FL. The reason why the one-shot FL performs worse than the multi-round FL in small models lies in the difference between the local model updates in Eq. 5 and Eq. 6. In Eq. 5, after the $t$-th communication round, the local training starts from the updated global model $\bm{w}^{(t,0)}$, which is aggregated by all the local models in $t$-th round and contains richer knowledge. Therefore, the client can compute a more accurate gradient $g_{i}(\bm{w}_{i}^{(t,j)})$ based on the updated model. On the contrary, in one-shot FL (Eq. 6), clients can only continuously train the local models without global information. The poor performance of one-shot FL is due to the gradients calculated on the local models being less accurate than those calculated on the aggregated global model. This local error can be expressed in mathematical form:
$\varepsilon_{i}=\sum_{j=k+1}^{Tk}\beta_{i}^{(0,j)}[(g_{i}(\bm{w}_{i}^{(0,j)})-g_{i}(\bm{w}_{i}^{(t,j-kt)})],\quad\text{where}\ t=\lceil\frac{j}{k}\rceil,$ (7)
where $\lceil\cdot\rceil$ means ceiling. $\varepsilon_{i}$ is the accumulated local update difference between one-shot FL and multi-round FL. Since the global model is aggregated by local models, the global model update difference $\varepsilon_{i}$ is the aggregation of local differences. Its L2 norm can then be bounded by the sum of local differences with triangle inequality, which is:
$\|\varepsilon\|\leq\sum_{i=1}^{m}\|\varepsilon_{i}\|.$ (8)
The global difference can be further simplified by the following assumptions.
@@
Assumption 2 (Bounded Model Updates). The L2 norm of model updates during FL are much smaller than the initial model parameters, that is, $\|\bm{w}^{(t,j)}-\bm{w}^{(0,0)}\|\leq\tau\|\bm{w}^{(0,0)}\|,\ 0<\tau<1$.
@@
Theorem 1 (The one-shot global difference is related to $L$, $\tau$, epoch numbers $Tk$, and number of clients $m$). Under Assumptions 1 and 2, ignoring the learning rates, the difference between one-shot FL and multi-round FL can be bounded as follows:
$\|\varepsilon\|\leq\Gamma\|\bm{w}^{(0,0)}\|,\ \text{where}\ \Gamma=L\tau Tkm.$ (9)
@@
This equation indicates that with lower values of $L$, $\tau$, $T$, $k$, and $m$, the model update of one-shot FL will be closer to that of multi-round FL. Since our experiments have shown that LLMs exhibit significant advantages over small models in one-shot learning, we conduct experiments on the factors in Equation 9 to provide a detailed explanation of this phenomenon.
Proof of Theorem 1 According to Eq. 7 and Eq. 8, ignoring the learning rates, the difference of the global model can be bounded by:
Considering Assumption 1, we can further factor out the effect of client heterogeneity and write the global model’s cumulative error in an $m\times\varepsilon_{i}$” form. Specifically, the aggregate deviation satisfies:
$\begin{split}\|\varepsilon\|\leq\sum_{j=k+1}^{Tk}Lm\|(\bm{w}_{i}^{(0,j)}-\bm{w}_{i}^{(t,j-kt)}\|,\end{split}$ (11)
According to Assumption 2, we can deduce:
$\begin{split}\|\varepsilon\|\leq\sum_{j=k+1}^{Tk}L\tau m\|\bm{w}^{(0,0)}\|,\end{split}$ (12)
Thus we have:
$\begin{split}\|\varepsilon\|\leq L\tau Tkm\|\bm{w}^{(0,0)}\|,\end{split}$ (13)
which is Theorem 1. This proof provides the intuition that, smoother model with smaller model updates and less training steps during FL has smaller one-shot global difference, thus performs better in one-shot FL. Motivated by this intuition, we use experimental evidence to show that FMs achieve better one-shot performance in the following paragraphs.
(a) The estimated $L$.
(b) The estimated $\tau$.
(c) The $\|\bm{w}^{(0,0)}\|$ values.
Figure 2: Experiment results on $L$, $\tau$, and $\|\bm{w}^{(0,0)}\|$ in different models. We use the CIFAR-10 dataset to compute the gradient on ResNet18 [he2016deep]. We use the WizardLM dataset [xu2023wizardlm] to compute the gradient on the language models. Models to the left of the red dashed line are small models, while those to the right are foundation models (FMs). The figures indicate that large FMs have significantly smaller $L$ and $\tau$ values compared to small models. Additionally, $\|\bm{w}^{(0,0)}\|$ does not increase proportionally with the model size. In conclusion, without considering other unrelated influencing factors, the value of $\Gamma\|\bm{w}^{(0,0)}\|$ decreases as the model size increases.